# MVP-Shot: Multi-Velocity Progressive-Alignment Framework for Few-Shot Action Recognition

### v1

@@ IV-D Ablation Study
Influence of different backbones. The previous comparison results are all based on CLIP-RN50. Here, we conduct detailed experiments to evaluate the impact of different backbones. We adopt the CLIP-ViT-B backbone, and the other settings are the same as those of our work. We list comparison results with CLIP-FSAR based on the CLIP-ViT-B backbone in Table VII. We conduct the experiments on on four gold-standard datasets (i.e., HMDB51 [19], UCF101 [20], kinetics100 [21], and SSv2-small [22]). From Table VII, we can see our MVP-shot based on CLIP-ViT-B consistently outperforms CLIP-FSAR [60], demonstrating the effectiveness of our method. In addition, MVP-shot based on CLIP-ViT-B generally achieves results superior to that based on CLIP-RN50, suggesting that a stronger pre-trained model contributes to enhanced few-shot generalization.
@@
Testing of $N$-Way Few-Shot Action Recognition. In the previous experiments, all of our comparative evaluation experiments are carried out under the $5$-way setting. To further explore the effect of varying $N$ on the few-shot performance, we compare $N$-way $1$-shot results on Kinetics in Fig. 4, where $N$ varies from $5$ to $10$. Results show as $N$ increases, the classification difficulty becomes higher. For example, the $10$-way result of MVP-shot decreases by $7.0$% compared to the $5$-way result. Nevertheless, the performance of our MVP-shot is still consistently ahead of the comparison methods, such as CLIP-FSAR [60]. This demonstrated that our proposed MVP-shot can boost performance by jointly exploring multi-velocity feature representation and multi-velocity feature alignment.
Figure 5: Comparison results (§IV-D) with different numbers of support samples in $5$-way $K$-shot setting on kinetics [21] dataset.
@@
$5$-way $K$-shot Few-shot Classification. To more comprehensively assess the performance of our model in various few-shot scenarios, we conduct extra experiments on Kinetics [21] by increasing the number of support samples from $1$-shot to $5$-shot. As shown in Fig. 5, as $K$ decreases, the classification difficulty becomes higher, and the performance decreases. Though too few support samples limit the performance of MVP-shot and CLIP-FSAR [60], MVP-shot performs consistently better than baseline and CLIP-FSAR in all settings. This demonstrates that our proposed MVP-shot can boost performance by exploring multi-velocity feature matching.
@@
Figure 6: Similarity visualization between query samples ($q_{n}$) and support prototypes ($s_{n}$) with different methods in a meta-test episode from Kinetics. The action classes, from left to right, are Shot Put, Catching or Throwing Baseball, Playing Ice Hockey, Drop Kicking and Dribbling Basketball, respectively. A higher score indicates a greater degree of similarity. The green box indicates correct prediction and the red box indicates incorrect prediction.
@@ IV-E Visualization Analysis
Similarity Visualization. To qualitatively demonstrate the proposed multi-velocity feature alignment in our method, we visualize the predicted similarities between query and support prototypes for an episode under both single-velocity and multi-velocity settings in Fig. 6. Compared to single-velocity alignment, multi-velocity alignment captures pair-wise semantic relevance at different velocity scales, compensating for the shortcomings of single-velocity alignment. Thus, our proposed multi-velocity alignment makes more accurate decisions. Specifically, for the fourth query sample in Kinetics, the incorrect decision obtained by the single-velocity (n=$1$) matching can be rectified by fusing the results of multi-velocity matching. Experimental results demonstrate that multi-velocity matching Complements each other, and combines alignment across all velocity features, achieving the best performance.
Attention Visualization of Our MVP-shot. Fig. 7 shows the attention visualization of our MVP-shot on UCF101 [20] and HMDB51 [19] in the $5$-way $1$-shot setting. according to the input RGB image sequences, the attention maps of the multi-modal transformer model using CLIP pre-trained weights, which we have mentioned in §IV-D, are compared to the attention maps with our MVP-shot, which refers to Progressive Semantic-Tailored Interaction module (PSTI). The attention map generated with multi-modal transformer often contains a significant number of unrelated or distracting focus areas. In contrast, the attention maps generated by our MVP-shot, which uses the PSTI module, focus more on action-related objects and reduce attention to the background and unrelated objects. Specifically, the frames in “Ride Horse” focus on the person on the horse, and the frames in “Ice Dancing” focus on dancers on the ice. These observations provide empirical evidence of the effectiveness of our PSTI module in enhancing semantic and spatiotemporal representation.
Figure 7: Attention visualization of our MVP-shot on HMDB51 [19] and UCF101 [20] in the 5-way 1-shot setting. Corresponding to the original RGB images (left), the attention maps of the multi-modal transformer model (middle) are compared to the attention maps of our MVP-shot (right). The temporal alignment metric is OTAM [11].
@@ V Conclusion
In this work, we first point out that single-scale feature alignment in previous FSAR methods is insufficient to reflect the speed diversity of action instances. To alleviate this problem, we propose a novel MVP-shot framework to obtain and align multi-velocity action features in a progressive way. A Progressive Semantic-Tailored Interaction module is designed to capture trustworthy multi-velocity feature representation. Subsequently, a Multi-Velocity Feature Alignment module is presented to measure the similarity of multi-velocity action features. By this means, our MVP-shot enables robust and comprehensive few-shot matching. Experimental results demonstrate that MVP-shot achieves state-of-the-art performance on the four standard benchmarks.

### v4

@@ IV-D Ablation Study
Influence of Different Backbones. The previous comparison results we reported are all based on CLIP-RN50 visual encoder. Here, we conduct detailed experiments to evaluate the influence of different backbones. We adopt the CLIP-ViT-B backbone, and the other settings are the same as those of our work. We list comparison results with CLIP-FSAR based on the CLIP-ViT-B backbone in Table VII. We conduct experiments on four widely used benchmarks (i.e., HMDB51 [19], UCF101 [20], kinetics100 [21], and SSv2-small [22]). From Table VII, we can see based on CLIP-ViT-B, our MVP-shot consistently outperforms CLIP-FSAR [60], demonstrating the superiority of our method. In addition, our MVP-shot model utilizing CLIP-ViT-B generally outperforms its counterpart based on CLIP-RN50, suggesting that a more robust pre-trained visual model leads to enhanced few-shot generalization.
@@
Testing of $N$-Way Few-Shot Action Recognition. We carry out all of the comparative evaluation experiments under the $5$-way setting in the previous experiments. To further investigate the effect of varying categories $N$ on the few-shot performance, we compare $N$-way $1$-shot results on Kinetics (as shown in Fig. 4). The results indicate that as the value of $N$ increases, the classification accuracy gradually becomes more challenging. For instance, the $10$-way performance of MVP-shot exhibits a decrease of $7.0$% compared to the 5-way scenario. Nevertheless, the performance of our MVP-shot is still consistently better than the comparison methods, such as CLIP-FSAR [60]. This demonstrated that our proposed MVP-shot can boost performance by jointly exploring multi-velocity feature representation and multi-velocity feature alignment.
@@
$5$-way $K$-shot Few-shot Classification. To more comprehensively assess our model’s performance across diverse few-shot scenarios, we conduct additional experiments by progressively increasing the number of support samples from $1$-shot to $5$-shot on Kinetics [21]. As $K$ decreases, the classification difficulty becomes higher, and the performance decreases (as shown in Fig. 5). Though the performance of MVP-shot and CLIP-FSAR [60] is limited due to too few support samples, MVP-shot performs consistently better than baseline and CLIP-FSAR in all settings. This demonstrates that our proposed MVP-shot can boost performance by exploring multi-velocity feature matching.
Figure 5: Comparison results (§IV-D) with different numbers of support samples on kinetics [21] dataset under $5$-way $K$-shot settings.
@@
Figure 6: Similarity visualization between support prototypes ($s_{n}$) and query samples ($q_{n}$) with different methods in an episode task sampled from Kinetics. The action classes, from left to right, are Shot Put, Catching or Throwing Baseball, Playing Ice Hockey, Drop Kicking and Dribbling Basketball, respectively. The correct prediction is indicated by the green box, while the red box denotes an incorrect prediction.
@@ IV-E Visualization Analysis
Similarity Visualization. To qualitatively demonstrate the proposed multi-velocity feature alignment in our MVP-shot, we visualize the predicted similarities between query videos and class prototypes for an episode under both single-velocity and multi-velocity settings in Fig. 6. Compared to single-velocity alignment, multi-velocity alignment captures pair-wise semantic relevance at different velocity scales, compensating for the shortcomings of single-velocity alignment. Thus, our proposed multi-velocity alignment makes more accurate decisions. Notably, for the fourth query sample from Kinetics, the erroneous decision made by the single-velocity (n=$1$) matching can be rectified by the fusion of multi-velocity matching results. Experimental results demonstrate that multi-velocity matching Complements each other, and combines alignment across all velocity features, achieving the best performance.
${}_{\!}$ ${}_{\!}$ ${}_{\!}$ Our ${}_{\!}$ We visualize the attention maps of our MVP-shot on HMDB51 [19] and UCF101 [20] under the $5$-way $1$-shot setting (as shown in Fig. 7). According to the input RGB image sequences, we compare the attention maps generated by the multi-modal transformer model using pre-trained CLIP model, as discussed in §IV-D, with the attention maps produced by our MVP-shot, which refers to Progressive Semantic-Tailored Interaction module (PSTI). The attention map generated with a multi-modal transformer often contains a significant number of unrelated or distracting focus areas. In contrast, the attention maps generated by our MVP-shot, which uses the PSTI module, reduce attention towards unrelated objects and the background and focus more on action-related objects. Specifically, the frames in “Ride Horse” focus on the person on the horse, and the frames in “Ice Dancing” focus on dancers on the ice. These attention Visualization maps provide empirical evidence of the efficacy of our PSTI module, which make use of semantic information provided by class names to focus on action-related areas.
Figure 7: Attention visualization of our MVP-shot on HMDB51 [19] and UCF101 [20] in the 5-way 1-shot setting. Corresponding to the original RGB images, the attention maps of the multi-modal transformer model are compared to the attention maps produced by our MVP-shot. We adopt OTAM [11] as the temporal alignment metric.
@@ V Conclusion
In this work, we first point out that single-scale feature alignment in previous FSAR methods is insufficient to reflect the speed diversity of action instances. To alleviate this problem, a novel MVP-shot framework is proposed to obtain and align multi-velocity action features in a progressive way. A Progressive Semantic-Tailored Interaction module is designed to capture trustworthy multi-velocity feature representation. Subsequently, a Multi-Velocity Feature Alignment module is presented to measure the similarity of multi-velocity action features. By this means, our MVP-shot enables more robust and more accurate few-shot feature matching. Experimental results demonstrate that MVP-shot yields solid performance gains and surpasses all the state-of-the-arts on the four standard benchmarks.